# Refining the WASP-132 multi-planetary system: discovery of a cold giant planet and mass measurement of a hot super-Earth

### v1

@@ 3.2.1 Rossiter McLaughlin Effect
The Rossiter-McLaughlin effect (Rossiter, 1924; McLaughlin, 1924) occurs during a planetary transit and impacts RVs from a typical Keplerian motion. We estimated the expected RV amplitude of the Rossiter-McLaughlin effect for both transiting planets with the classical method (Eq. 40 from Winn 2010) using our expected v and stellar radius measurements along with our derived values of the planets’ radii and impact parameters, which we later describe in Section 3.3. We find Rossiter-McLaughlin effect RV amplitudes of 1.55 m s-1 for planet c and 47.36 m s-1 for planet b. Thus it is necessary to remove RV data that occurred during transits to not affect the RV model parameters. Using the period and transit duration that we later describe in Section 3.3, we find three CORALIE RVs and four HARPS RVs that occur during a transit of planet c. We removed the four HARPS RVs from our analysis, but given the relatively small impact the planet c Rossiter-McLaughlin effect will have on CORALIE data we did not remove the three CORALIE RVs. We also find one CORALIE RV and five HARPS RVs that are during a transit of planet b and we remove all of these data from our analysis. The RVs that occur during a planet transit are noted in Appendix Tables A.1, A.2, and A.3. We note that high resolution ESPRESSO (Pepe et al., 2021) observations of the Rossiter-McLaughlin signal for both planets were obtained and are in preparation for publication by a separate team.
3.2.2 HARPS two-planet residuals analysis
@@
Our CORALIE and HARPS RVs clearly show the signal of the 7-day hot Jupiter and a longer period companion around 1800 days as seen in the generalized Lomb-Scargle (GLS) periodogram (Lomb, 1976; Scargle, 1982; Zechmeister & Kürster, 2009) implemented in the astropy (Astropy Collaboration et al., 2013, 2018) Python package of all the WASP-132 RVs in Figure 1. The astropy periodogram package can calculate False Alarm Probabilities (FAP) using various methods and we use the default Baluev (2008) method. Notably, we see the 7-day and long-period companion as well as one-day signals above the 0.1% FAP level in the periodogram. We also detected that a long-term RV trend is within the data.
@@
In order to detect the much smaller signal from the 1-day period super-Earth planet we first fit the longer period signals and analyzed the residuals. For our initial RV analysis, we used the software suite pyaneti444https://github.com/oscaribv/pyaneti (Barragán et al., 2019, 2022), which couples a Bayesian framework with a Markov chain Monte Carlo (MCMC) sampling to produce posterior distributions of the fitted parameters. We first fit the RVs with two Keplerians for the hot Jupiter and long-period outer companion as well as a linear trend. We note that we fit the data without a linear trend and the linear trend model is preferred with a difference in Bayesian Information Criterion (BIC) of 134. We put Gaussian priors on the period of the hot Jupiter based on our transit analysis presented in Section 3.3. With our initial two-planet fit we find the outer companion to have a period of 1792 days and RV semi-amplitude = 105.6 m s-1 and a linear trend of 23.5 m s-1 year-1. After removing these signals we analyzed the HARPS RVs which have the necessary precision to detect the smaller 1-day planet.
@@
Noticeably the HARPS RV residuals of the 2-planet plus trend fit have variation much larger than the expected signal from the 1-day period super-Earth planet and we find strong signals at 21 and 10 days as displayed in Figure 2. It is well known that magnetically active regions (spots and plages) can induce periodic and quasi-periodic Doppler signals at the stellar rotation frequency and its harmonics (e.g., Boisse et al., 2011). Our log(R’HK) = 4.852 0.039 value suggests that the star may be magnetically active and additionaly Hellier et al. (2017) found a possible rotational modulation with a period of 33 3 days and an amplitude of 0.4 mmag in three out of five seasons of WASP data. We expect these signals in the HARPS data are due to magnetic activity and rotation of the star. We also see repeated signals in the in the various HARPS activity indicators including peak periods of 28 days in the Cross Correlation Function (CCF) Full-Width at Half-Maximum (FWHM), 10 days in the CCF Bisector Span (BIS), and 30 in the Mt Wilson ‘S index’ SMW, Na lines, and Ca lines. We show the GLS periodograms for the CCF FWHM, CCF BIS, and SMW in Figure 2.
@@
In order to account for this activity signal we applied a multi-dimensional Gaussian process (GP) approach, the pyaneti implementation of which is as described in Rajpaul et al. (2015). We use the quasi-periodic (QP) kernel, as defined by Roberts et al. (2013) and described in Barragán et al. (2022), and place an uninformative prior with a range of 5 to 45 days for the period of the QP kernel. We tested combinations of available activity indicators and in all cases the period of QP kernel was found to be at PGP 30-32 days, which we take to be the rotation period of the star. We found that a two-dimensional GP fit with the RV and BIS activity indicators gave the best results and obtained the lowest BIC value and smaller HARPS RV jitter value when we allow an RV jitter to be fit. With this fit we find a RV semi-amplitude of = 4.61 m s-1 corresponding to a mass of Mp = 6.2 M⊙ for the 1-day planet and a PGP = 31.22 days. We show the multi-dimensional GP fit with BIS and the HARPS RV residuals in Figure 3.
@@
For the previous fits we fixed the eccentricity of the 1-day period planet to 0. We also tested for eccentricity by fitting for cos and sin. The fit finds an eccentricity of e = 0.20. However, the BIC is actually smaller by 3 for the fit with the eccentricity. Given the large errors on the eccentricity, the smaller BIC of the fixed eccentricity, and the physical likelihood that the planet is tidally locked we fix the eccentricity to 0. Additionally given the complexity of the activity and small planet signal we conclude we do not have the precision to measure the eccentricity of the 1-day planet with our HARPS data.
@@
Given the short period of the 1-day planet signal and longer rotation period of the star, we also tested fitting only HARPS RVs that had multiple observations taken on a given night and fit each night as its own offset in order to constrain the RV semi-amplitude of only the planet signal, similar to the analysis presented in Hatzes et al. (2011) to obtain the mass of CoRoT-7b (Léger et al., 2009; Queloz et al., 2009). After removing RVs during transits, we had 12 different nights of HARPS data with two observations per night. Allowing each night to have its own RV offset and then fitting for a single Keplerian we found an RV semi-amplitude of = 5.03 m s-1 corresponding to a mass of Mp = 6.7 M⊙, which is consistent with our Keplerian plus GP model of the residuals. However, the orbital phase coverage is not optimal which may explain the small differences with other analyses.
@@
3.2.3 Three-planet RV-only Model
@@
Finally for our RV-only analysis we perform a three-planet fit plus a linear trend and a QP GP kernel. We only use a 1-dimensional GP as pyaneti cannot fit activity indicators for one instrument and not others, and we do not find a good fit when including the CORALIE activity indicators which are essentially noise when trying to fit the activity signal. We put Gaussian priors on the periods and ephemerides of the two inner planets based on our transit fit in Section 3.3. For the outer planet period we used uniform priors between 1700 and 1900 days. We also put uniform priors on the QP GP period between 30 and 32 days based on our results from the 2-dimensional GP modeling with the HARPS residuals. We fixed the 1-day planet’s eccentricity to 0. All other parameters were given uniform priors. We show our full fit to the RV data in Figure 5 and present the fit parameters including our mass measurements in Table 2, which is in agreement with our previous analyses.
@@ 3.3 Joint Photometry and RV analysis
We modeled both the photometry and RV data with juliet (Espinoza et al., 2019), which uses Bayesian inference to model a set number of planetary signals using batman (Kreidberg, 2015) to model the planetary transits and radvel (Fulton et al., 2018) to model the RVs. Of note, although juliet cannot handle multi-dimensional Gaussian processes (GPs) for RV data, for light curves juliet can model stellar activity as well as instrumental systematics with GPs (e.g., Gibson, 2014) or simpler parametric functions, which we do not find available in pyaneti. For the transit model, juliet performs an efficient parameterization by fitting for the parameters and to ensure uniform exploration of the (planet-to-star ratio; ) and (impact parameter) parameter space. We used the nested sampling method dynesty (Speagle, 2019) implemented in juliet with 1500 live points and we ran the fit until the estimated uncertainty on the log-evidence was smaller than 0.1.
We added a white-noise jitter term in quadrature to the error bars of both the photometry and RV data to account for underestimated uncertainties and additional noise that was not captured by the model. To account for possible residual systematics and activity affecting the transit fit of TESS Sectors 11, 38, and 65, we fit a Gaussian process (GP) using a Matérn-3/2 kernel via celerite (Foreman-Mackey et al., 2017) within the juliet framework. In addition to the stellar activity fitting presented in Section 3.2.2, a GP can account for correlated noise of various origins and propagates the uncertainty (e.g., Gibson et al., 2012; Gibson, 2014). Additionally, properly fitting out of transit data will more accurately set the baseline for in transit data. We display the GP fit to the full TESS light curves in Figure 14. We also fit a Gaussian process (GP) using a Matérn-3/2 kernel for the RV data and put uniform priors on the time-scale based on our initial RV analysis in Section 3.2.3. We fit the CHEOPS data by detrending with the background flux and the roll angle by using cos(), cos(), sin(), and sin(2). We display our full fit to the CHEOPS data in Figure 15.
@@
Period (days) | |||
Time of transit center (BJD) | |||
@@
Time of transit center (BJD) | |||
@@
* indicates a uniform distribution between and ; a Jeffrey or log-uniform distribution between and ; a normal distribution with mean and standard deviation ; and a Beta prior as detailed in Kipping (2014). Parameters with no prior distribution were derived. We sample from a normal distribution for the stellar mass, stellar radius, and stellar temperature, that are based on the results from Section 3.1 to derive parameters.
@@
For limb darkening, we derived quadratic coefficients and their uncertainties for different photometric filters using the LDCU555https://github.com/delinea/LDCU routine (Deline et al., 2022) and used them to set Gaussian priors on the limb-darkening parameters. We used uniform priors for the period and transit center times. We fixed the eccentricity for the 1-day period planet and added a Beta prior for transiting planets for the hot Jupiter, as described in Kipping (2014). We used the (stellar density) as a parameter instead of the scaled semi-major axis (/R⋆). The normal prior on stellar density is informed by the stellar analysis in Section 3.1.
For our final RV and photometry joint fit we constrained the RV semi-amplitude of planet c as well as the period, , eccentricity, RV , and of planet d using values from our initial pyaneti RV analysis presented in Section 3.2.2. We tested several times putting less constraints on these parameters within the juliet global model but found that the model would get stuck with solutions that had lower likelihoods than our final solution. We therefore used priors from our initial pyaneti analysis to guide the juliet global model to find the most likely solution. We account for these priors in our uncertainties of the modeled and derived juliet parameters.

### v2

@@ 3.2.1 Rossiter McLaughlin effect
The Rossiter-McLaughlin effect (Rossiter, 1924; McLaughlin, 1924) occurs during a planetary transit and impacts RVs from a typical Keplerian motion. We estimated the expected RV amplitude of the Rossiter-McLaughlin effect for both transiting planets with the classical method (Eq. 40 from Winn 2010) using our v and stellar radius measurements along with our derived values of the planets’ radii and impact parameters, which we later describe in Sect. 3.4. We find Rossiter-McLaughlin effect RV amplitudes of 1.6 m s-1 for planet c and 47 m s-1 for planet b, assuming = 0 (no obliquity). Thus it is necessary to remove RV data that occurred during transits to not affect the RV model parameters. Using the period and transit duration that we later describe in Sect. 3.4, we find three CORALIE RVs and four HARPS RVs that occur during a transit of planet c. We removed the four HARPS RVs from our analysis, but given the relatively small impact the planet c Rossiter-McLaughlin effect will have on CORALIE data, due to the CORALIE RV uncertainties, we did not remove the three CORALIE RVs. We also find one CORALIE RV and five HARPS RVs that are during a transit of planet b and we remove all of these data from our analysis. We note that high resolution ESPRESSO (Pepe et al., 2021) observations of the Rossiter-McLaughlin signal for both planets were obtained and are in preparation for publication by a separate team.
3.2.2 HARPS linear trend plus two-planet residuals analysis
@@
Our CORALIE and HARPS RVs clearly show the signal of the 7-day hot Jupiter and a longer period companion around 1800 days as seen in the generalized Lomb-Scargle (GLS) periodogram (Lomb, 1976; Scargle, 1982; Zechmeister & Kürster, 2009) implemented in the astropy (Astropy Collaboration et al., 2013, 2018) Python package of all the WASP-132 RVs in Fig. 1. The astropy periodogram package can calculate False Alarm Probabilities (FAP) using various methods and we use the default Baluev (2008) method. We accounted for RV offsets between the instruments using the best-fitting systemic velocities, as derived from the COR07, COR14, and HARPS datasets (see Table 2). Notably, we see the 7-day and long-period companion as well as one-day signals above the 0.1% FAP level in the periodogram. We also detected that a long-term RV trend is within the data.
@@
In order to detect the much smaller signal from the 1-day period super-Earth planet we first fit the longer period signals and analyzed the residuals. For our initial RV analysis, we used the software suite pyaneti444https://github.com/oscaribv/pyaneti (Barragán et al., 2019, 2022), which couples a Bayesian framework with a Markov chain Monte Carlo (MCMC) sampling to produce posterior distributions of the fitted parameters. We first fit the RVs with two Keplerians for the hot Jupiter and long-period outer companion as well as a linear trend. We note that we fit the data without a linear trend and the linear trend model is preferred with a difference in Bayesian Information Criterion (BIC) of 134. We put Gaussian priors on the period of the hot Jupiter based on our transit analysis presented in Sect. 3.4. With our initial two-planet fit we find the outer companion to have a period of 1792 days and RV semi-amplitude = 105.6 7.4 m s-1 and a linear trend of 23.5 m s-1 year-1. After removing these signals we analyzed the HARPS RVs which have the necessary precision to detect the smaller 1-day planet.
@@
Noticeably the HARPS RV residuals of the 2-planet plus trend fit have variation much larger than the expected signal from the 1-day period super-Earth planet and we find strong signals at 21-28 and 8-11 days as displayed in Fig. 2. It is well known that magnetically active regions (spots and plages) can induce periodic and quasi-periodic Doppler signals at the stellar rotation frequency and its harmonics (e.g., Boisse et al., 2011). Our log(R’HK) = 4.852 0.039 value suggests that the star may be magnetically active and additionally Hellier et al. (2017) found a possible rotational modulation with a period of 33 3 days and an amplitude of 0.4 mmag in three out of five seasons of WASP data. We expect these signals in the HARPS data are due to magnetic activity and rotation of the star. We also detect significant signals in the various HARPS activity indicators. We display GLS periodograms in the period domain of the CCF FWHM, CCF BIS, and Mt Wilson S index SMW in Fig. 2. We present a more detailed view of the HARPS activity indicator data in Sect. 3.3.
@@
In order to account for this activity signal we applied a multidimensional Gaussian process (GP) approach, the pyaneti implementation of which is as described in Rajpaul et al. (2015). We used the quasi-periodic (QP) kernel, as defined by Roberts et al. (2013) and described in Barragán et al. (2022), and placed an uninformative prior with a range of 5 to 45 days for the period of the QP kernel. We tested combinations of available activity indicators and in all cases the period of QP kernel was found to be at PGP 30-32 days, which we take to be the rotation period of the star. We found that a two-dimensional GP fit with the RV and BIS activity indicators gave the best results and obtained the lowest BIC value and smaller HARPS RV jitter value when we allow an RV jitter to be fit. With this fit we found a RV semi-amplitude of = 4.61 m s-1 corresponding to a mass of Mp = 6.2 M⊕ for the 1-day planet and a PGP = 31.22 days. We show the multidimensional GP fit with BIS and the HARPS RV residuals in Fig. 3.
@@
For the previous fits we fixed the eccentricity of the 1-day period planet to 0. We also tested for eccentricity by fitting for cos and sin. The fit finds an eccentricity of e = 0.20. However, the BIC is actually higher by 3 for the fit with the eccentricity. Given the large errors on the eccentricity, the smaller BIC of the fixed eccentricity, and the physical likelihood that the planet is tidally locked we fix the eccentricity to 0. Additionally given the complexity of the activity and small planet signal we conclude we that we cannot measure the eccentricity of the 1-day planet with the current RV dataset.
@@
Given the short period of the 1-day planet signal and longer rotation period of the star, we also tested fitting only HARPS RVs that had multiple observations taken on a given night and fit each night as its own offset in order to constrain the RV semi-amplitude of only the planet signal, similar to the analysis presented in Hatzes et al. (2011) to obtain the mass of CoRoT-7b (Léger et al., 2009; Queloz et al., 2009). After removing RVs during transits, we had 12 different nights of HARPS data with two observations per night. Allowing each night to have its own RV offset and then fitting for a single Keplerian we found an RV semi-amplitude of = 5.03 m s-1 corresponding to a mass of Mp = 6.7 M⊕, which is consistent with our Keplerian plus GP model of the residuals.
@@
3.2.3 Linear trend plus three-planet RV-only model
@@
Finally for our RV-only analysis we perform a three-planet fit plus a linear trend and a QP GP kernel. We only use a one-dimensional GP as pyaneti cannot fit activity indicators for one instrument and not others, and we do not find a good fit when including the CORALIE activity indicators which are essentially noise when trying to fit the activity signal. We put Gaussian priors on the periods and ephemerides of the two inner planets based on our transit fit in Sect. 3.4. For the outer planet period we used uniform priors between 1700 and 1900 days. We also put uniform priors on the QP GP period between 30 and 32 days based on our results from the two-dimensional GP modeling with the HARPS residuals. We fixed the 1-day planet’s eccentricity to 0. All other parameters were given uniform priors. We show our full fit to the RV data in Fig. 5 and present the fit parameters including our mass measurements in Table 2, which is in agreement with our previous analyses. Table 2 presents the QP GP hyper-parameters including A0 and A1, which are the coefficients of the GP and its derivative, respectively. A0 and A1 are amplitudes that act as scale factors, determining the typical deviation from the mean function. is the long-term evolution time-scale (the lifetime of active regions). is the inverse of the harmonic complexity (how complex variations are inside each period, which is related to distribution of active regions). is the long-term evolution time-scale (the lifetime of active regions). In Table 2 T0d refers to the time of inferior conjunction.
3.3 RV signal and activity analysis
Here we further investigate the nature of the signals in the CORALIE and HARPS RVs. We employ a form of the pre-whitening procedure (e.g., Queloz et al., 2009; Hatzes, 2013; Perger et al., 2017) based on a Fourier transform (FT) frequency analysis. In this procedure, we subtract RV signals, removing their corresponding FT peaks from the data, and then analyze the resulting generalized Lomb-Scargle (GLS) periodograms in the frequency domain. We also created spectral window functions (Roberts et al., 1987; Dawson & Fabrycky, 2010), the Fourier transform of the observation times, for the CORALIE and HARPS RVs. The window function can help identify aliases in time series data as detailed by Dawson & Fabrycky (2010).
In the top panel of Fig. 3 of the Supplementary Material published on Zenodo, we present a GLS periodogram in the frequency domain for the window function of the CORALIE RVs. The second panel displays the GLS periodogram of the CORALIE RVs. We account for the RV offsets between COR07 and COR14. The third through sixth panels display the GLS periodograms of the CORALIE RVs subtracted by the long-term linear trend; the linear trend and the Doppler signal of planet b; the linear trend and the Doppler signal of planet d; and the linear trend and the Doppler signals of planet d and b; respectively. From this analysis we find a strong signal at planet d’s corresponding frequency after removing the linear trend and planet b’s Doppler signal. We also see the FT peaks around 1 day-1 are removed after subtracting the linear trend and planet d.
In addition, we examined GLS periodograms of the CORALIE activity indicators. The top panel of Fig. 4 of the Supplementary Material published on Zenodo again displays the GLS periodogram for the window function of the CORALIE data. Panels two through seven display display GLS periodograms of the CCF FWHM, CCF BIS, CCF Contrast, and the chromospheric indices measured in H, Na, and Ca ii H&K lines lines, respectively. The offsets between the COR07 and COR13 activity indicators are negligible. From the CORALIE activity indicators we do not examine any significant peaks around planet d’s frequency that would suggest it originated from stellar activity. We also do not see any long-term trend originating from a magnetic cycle that could explain the linear RV trend.
We performed similar procedures with the HARPS data as displayed in Figs. 5 and 6 of the Supplementary Material published on Zenodo. Figure 5 of the Supplementary Material published on Zenodo displays GLS periodograms in the frequency domain of the HARPS window function, HARPS RVs, and the RVs with various signals subtracted including the linear trend and Doppler signals of planets d, b, and c. For the activity signal we fit the phase and amplitude of sinusoidal signals at frequencies of the approximate stellar rotation period of 31.4 days and its first two harmonics of 15.7 and 10.5 days. Although the signal is not significant, we see a small peak around planet c’s frequency after subtracting the other signals and the activity signals. The GLS periodograms in the frequency domain of the HARPS activity indicators are displayed in Fig. 6 of the Supplementary Material published on Zenodo which includes the Mt Wilson S index SMW. We detect significant signals around the stellar rotation frequency of 1/31.4 days in the CCF FWHM, CCF Contrast, H, Na, Ca ii, and SMW GLS periodograms.
@@ 3.4 Joint photometry and RV analysis
We modeled both the photometry and RV data with juliet (Espinoza et al., 2019), which uses Bayesian inference to model a set number of planetary signals using batman (Kreidberg, 2015) to model the planetary transits and radvel (Fulton et al., 2018) to model the RVs. Of note, although juliet cannot handle multidimensional Gaussian processes (GPs) for RV data, for light curves juliet can model stellar activity as well as instrumental systematics with GPs (e.g., Gibson, 2014) or simpler parametric functions, which we do not find available in pyaneti. For the transit model, juliet performs an efficient parameterization by fitting for the parameters and to ensure uniform exploration of the (planet-to-star ratio; ) and (impact parameter) parameter space. Espinoza (2018) details the and parameterization and the transformation to the and plane. We used the nested sampling method dynesty (Speagle, 2019) implemented in juliet with 1500 live points and we ran the fit until the estimated uncertainty on the log-evidence was smaller than 0.1.
We added a white-noise jitter term in quadrature to the error bars of both the photometry and RV data to account for underestimated uncertainties and additional noise that was not captured by the model. To account for possible residual systematics and activity affecting the transit fit of TESS Sectors 11, 38, and 65, we fit a Gaussian process (GP) using a Matérn-3/2 kernel via celerite (Foreman-Mackey et al., 2017) within the juliet framework. In addition to the stellar activity fitting presented in Sect. 3.2.2, a GP can account for correlated noise of various origins and propagates the uncertainty (e.g., Gibson et al., 2012; Gibson, 2014). Additionally, properly fitting out of transit data will more accurately set the baseline for in transit data. We display the GP fit to the full TESS light curves in Fig. 1 of the Supplementary Material published on Zenodo. We also fit a Gaussian process (GP) using a Matérn-3/2 kernel for the RV data and put uniform priors on the time-scale based on our initial RV analysis in Sect. 3.2.3. We fit the CHEOPS data by detrending with the background flux and the roll angle by using cos(), cos(), sin(), and sin(2). We display our full fit to the CHEOPS data in Fig. 2 of the Supplementary Material published on Zenodo.
@@
| Period (days) | (7.13350.1) | ||
| Time of transit center (BJD) | (2459337.60.2) | ||
@@
| Time of inferior conjunction (BJD) | |||
@@
* indicates a uniform distribution between and ; a Jeffrey or log-uniform distribution between and ; a normal distribution with mean and standard deviation ; and a Beta prior as detailed in Kipping (2014). Parameters with no prior distribution were derived. We sample from a normal distribution for the stellar mass, stellar radius, and stellar temperature, that are based on the results from Sect. 3.1 to derive parameters.
@@
For limb darkening, we derived quadratic coefficients and their uncertainties for different photometric filters using the LDCU555https://github.com/delinea/LDCU routine (Deline et al., 2022) and used them to set Gaussian priors on the limb-darkening parameters. We used uniform priors for the period and transit center times. We fixed the eccentricity for the 1-day period planet and added a Beta prior for transiting planets for the hot Jupiter, as described in Kipping (2014). We used the (stellar density) as a parameter instead of the scaled semi-major axis (/R⋆). The normal prior on stellar density is informed by the stellar analysis in Sect. 3.1.
For our final RV and photometry joint fit, we constrained the RV semi-amplitude of planet c as well as the period, , eccentricity, RV , and of planet d using values from our initial pyaneti RV analysis presented in Sect. 3.2.2. We tested several times putting fewer constraints on these parameters within the juliet global model but found that the model would get stuck with solutions that had lower likelihoods than our final solution. We therefore used priors from our initial pyaneti analysis to guide the juliet global model to find the most likely solution. We account for these priors in our uncertainties of the modeled and derived juliet parameters.